OpenStudy (roberts.spurs19):

Steady Flow Energy Equation help please? Steam enters an adiabatic nozzle at a speed of 20 m s-1, a temperature of 600C and pressure of 10 bar absolute. It exits at 400C and 1 bar absolute. (i) What is the exit speed?

OpenStudy (michele_laino):

we have to apply the Theorem of Bernoulli for gases. So we can write: \[\Large \frac{{v_2^2 - v_1^2}}{2} = \frac{{{I_1} - {I_2}}}{A}\] where: \[\Large \begin{gathered} {I_1} - {I_2} = {\text{difference of enthalpy}} \hfill \\ A = {\text{cross sectional area of the nozzle}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

the relationship for enthalpy is: \[\Large I = U + A \times p \times v\] where: \[\Large \begin{gathered} v = {\text{specific volume = volume/mass of gas}} \hfill \\ U = {\text{internal energy/mass of gas}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

p is the pressure of the gas